OpenStudy (anonymous):

solve the equation graphically by converting it to an equivalent equation with 0 on the right. 2x-5=√(x+4) am stuck

OpenStudy (whpalmer4):

To solve \(2x -5 = \sqrt{x+4}\) I would start by squaring both sides of the equation. Then I would move everything to one side so the other side = 0.

OpenStudy (anonymous):

would that take away the square root over the x+4 when I square it

OpenStudy (anonymous):

yes, you get \[(2x-5)^2=x+4\] or \[4x^2-20x+25=x+4\]